OpenStudy (hba):

Equation of a circle with center (-3,5) and radius 7 is :

OpenStudy (hba):

@waterineyes

OpenStudy (nurali):

the standard equation of a circle: (x-h)^2 + (y-k)^2 = r^2 given r=7 c(5,-3) (x-5)^2+(y-(-3))^2=7^2 (x-5)^2+(y+3)^2=49 answer

OpenStudy (hba):

@waterineyes I got x^2+y^2+6x-10y-15=0

OpenStudy (anonymous):

What is center??

OpenStudy (hba):

Note that i used the formula (x-a)^2+(y-b)^2=r^2 @waterineyes

OpenStudy (anonymous):

Center is (-3,5) and you are using (5,-3) Any special reason??

OpenStudy (hba):

The center (a,b)=(-3,5) and r=7

OpenStudy (anonymous):

Forget the formula.. Check once again for your solution..

OpenStudy (hba):

I am right :D

OpenStudy (anonymous):

Shouldn't it be : \((x+3)^2 + (y-5)^2 = 7^2\) ??

OpenStudy (hba):

I am not using Nuralis Formula Dude i have solved it right :D

OpenStudy (hba):

@waterineyes And I got x^2+y^2+6x-10y-15=0

OpenStudy (anonymous):

Explain to me what are you doing..

OpenStudy (hba):

@waterineyes (x+3)^2+(y-5)^2=7^2 x^2+6x+9+y^2-10y+25=49 x^2+y^2+6x-10y-15=0

OpenStudy (anonymous):

Oh my God.. Sorry, I was thinking that, that was your post.. Really sorry..

OpenStudy (anonymous):

That was of Nurali..

OpenStudy (hba):

Yeah lol :D Thanks anyway :)

OpenStudy (anonymous):

Yes you are right..